P= 2(a+b), then b=
Answer by Cromlix(4381)

Hi there,
P= 2(a+b)
Divide both sides by 2
P/2 = 2(a + b)/2
Cancel out the 2's on the right hand side
P/2 = (a + b)
Remove brackets on right hand side
P/2 = a + b
Subtract a from both sides
P/2 - a = a + b - a
P/2 - a = b
b = P/2 - a
Hope this helps :-)